# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 4.3. Positive stakes at Nash equilibrium
Derivation. Under $\mathbb{P}^{i}_{S_{-},S_{+}}$, let $\sigma_{i+1}\in\mathbb{N}_{+}\cup\{\infty\}$ denote the stopping time $\inf\big\{\ell\in\mathbb{N}_{+}:X_{\ell}=i+1\big\}$. In the specification (4) of Maxine’s net receipt $P_{+}$ as $T_{+}-C_{+}$, the running cost $C_{+}$ may be written $C_{+}\llbracket 1,t\rrbracket$ and $C_{+}\llbracket t+1,\infty\rrbracket$ where Maxine’s stakes up to the $t$th turn enter as summands in the first term. Using the strong Markov property at time $\sigma_{i+1}$, and dropping $C_{+}\llbracket 1,t\rrbracket\geq 0$, we obtain
@@
Here, the left-hand side equals $m_{i}$ by definition while right-hand side is
@@
The third right-hand term vanishes by Proposition 4.1, so that $m_{i}$ is bounded above by a weighted average of $m_{-\infty}$ and $m_{i+1}$. We will find as desired that $m_{i}\leq m_{i+1}$ by showing $m_{-\infty}\leq m_{i+1}$. In this regard, we first claim that $\mathbb{E}^{i+1}_{S_{-},0}[P_{+}]=m_{-\infty}$. To check this, note that Lemma 4.7(3) implies that that $S_{-}$ is wide. We may now make use of Proposition 4.8 to learn that $E_{-}$, and thus also $T_{+}=m_{-\infty}$, are $\mathbb{P}^{i}_{S_{-},0}$-almost certain. The absence of running costs for Maxine means that $P_{+}=T_{+}$ under $\mathbb{P}^{i+1}_{S_{-},0}$. The claim obtained, we use it and $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$ to find that $m_{i+1}\,=\,\mathbb{E}^{i+1}_{S_{-},S_{+}}[P_{+}]\,\geq\,\mathbb{E}^{i+1}_{S_{-},0}[P_{+}]=m_{-\infty}$, thereby confirming $m_{i}\leq m_{i+1}$. Omitting the similar proof that $n_{i+1}\leq n_{i}$, we obtain Lemma 4.11. ∎
@@
$m_{i}\,=\,\Big(\kappa\tfrac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i+1}+\Big(\kappa\tfrac{b_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i-1}\,-\,a_{i}\,.$ (33)
Proof. Maxine will spend $a_{i}$ at the first turn; the move will be stake with probability $\kappa$ and then she win it with conditional probability $\tfrac{a_{i}}{a_{i}+b_{i}}$; if she does so, the counter will reach $i+1$, and her resulting conditional mean receipt will be $m_{i+1}$; and this circumstance will equally arise if a fair coin lands heads on a flip move, with probability $(1-\kappa)/2$. Otherwise, Maxine’s receipt will be $m_{i-1}$. Note that the two ratios on the right-hand side of (33) are well defined, because $a_{i}+b_{i}>0$. ∎
@@
Proof of Proposition 4.10. By Lemma 4.7(3), $S_{-}$ is wide. By Lemma 4.14, $a_{i}>0$ implies that $a_{i+1}>0$. Hence, all coefficients $a_{i}$ are positive; by Lemma 4.14(2), so are all the $b_{i}$. By Lemma 4.13, the differences $m_{i,i+1}$ and $n_{i+1,i}$ are also found to be positive.∎
@@ 4.4. The forward implication
We are ready for the next derivation. The argument follows the lines of the proof of [26, Theorem $2.6(1)$], with a different approach used at the end to handle flip moves.
@@
$A_{+}-O_{+}\,=\,\bigg(\frac{\rho\,a_{i}^{\rho-1}b_{i}^{\rho}}{(a_{i}^{\rho}+b_{i}^{\rho})^{2}}\kappa\,m_{i-1,i+1}-1\bigg)\cdot\eta\cdot\big(1+o(1)\big)\,$, (34)
@@
$\rho\,a_{i}^{\rho}b_{i}^{\rho-1}\,\kappa\,n_{i+1,i-1}\;=\;\big(a_{i}^{\rho}+b_{i}^{\rho}\big)^{2}\,.$ (36)
The obtained equations (35) and (36) are ABMN(3,4)(3,4) with index $i$.
We have shown that $\big\{(a_{i},b_{i},m_{i},n_{i}):i\in\mathbb{Z}\big\}$ is an element of ABMN($\kappa,\rho$). To finish the proof of Theorem 1.8(1), it remains to confirm that the boundary values (6) are achieved. We will prove that $\lim_{i\to\infty}m_{-i}=m_{-\infty}$; the three other limits are similarly obtained. The sequence $\big\{m_{-i}:i\in\mathbb{N}\big\}$ decreases by Proposition 4.10 to a limit that we call $\mathfrak{m}_{-\infty}$.
By Definition 1.6 and $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$, $m_{i}=\mathbb{P}^{i}_{S_{-},S_{+}}[P_{+}]\geq\mathbb{P}^{i}_{S_{-},0}[P_{+}]$. It is Proposition 4.8 that now permits us to identify the right-hand term as being equal to $m_{-\infty}$. Hence, $\mathfrak{m}_{-\infty}\geq m_{-\infty}$; we wish to obtain the opposing inequality. We take the mean of the equality $P_{+}=T_{+}-C_{+}$ in (4) and remove non-negative running costs $C_{+}$ to find that $m_{i}\leq\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})\cdot m_{-\infty}+\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})\cdot m_{\infty}$ where we invoked Proposition 4.1 to eliminate a non-escape $E^{c}$ term. Thus $\mathfrak{m}_{-\infty}\leq m_{-\infty}$ provided that we show that $\lim_{i\to-\infty}\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})$ equals zero: far to the left, Mina’s victory is close to assured.
@@ 4.5. The reverse implication
Here we prove Theorem 1.8(2), the step at which the infinite-turn game is controlled by comparison with finite-trail counterparts. Throughout, $\big\{(a_{i},b_{i},m_{i},n_{i}):i\in\mathbb{Z}\big\}$ denotes an element of ABMN($\kappa,\rho$), with boundary data $(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})$ that satisfies (7). We define strategies $S_{-},S_{+}\in\mathcal{S}$ that offer $b$- and $a$-stake compatibly with the rule (5).
Since all counter moves are $\pm 1$, counter location is constrained by parity. First we denote the set of space-time sites that are thus in principle accessible for gameplay $X:\mathbb{N}\to\mathbb{Z}$ under $\mathbb{P}^{i}_{S_{1},S_{2}}$ for some strategy pair $(S_{1},S_{2})\in\mathcal{S}^{2}$.
@@
For $i\in\mathbb{Z}$, the forward play-cone $F_{i}$ of $i$ is set equal to
$F_{i}\,=\,\Big\{\,(k,\ell)\in\mathbb{Z}\times\mathbb{N}_{+}:|k-i|\leq\ell\,,\,|k-i|+\ell\in 2\mathbb{N}\,\Big\}\,.$
Let $S\in\mathcal{S}$ (and recall the formulation of the strategy space $\mathcal{S}$ from Section 1.3). A Mina deviation point is an element $(q,\ell)\in F_{i}$ for which there exists a trajectory $\psi:\llbracket 0,\ell\rrbracket\to\mathbb{Z}$ with $\psi(0)=i$ and $\psi(\ell)=q$ such that $S(\psi)\neq b_{q}$. Write $\mathsf{D}_{-}(S,i)\subseteq F_{i}$ for the set of Mina deviation points. The strategy $S$ is deviating for Mina if $\mathsf{D}_{-}(S,i)\neq\emptyset$. A Maxine deviation point is an element $(q,\ell)\in F_{i}$ such that $S(\psi)\not=a_{q}$ for some path $\psi$ as above. Write $\mathsf{D}_{+}(S,i)$ for the set of these points; if $\mathsf{D}_{+}(S,i)\not$$\emptyset$, then $S$ is deviating for Maxine.
@@
Write $P^{j,k}$ for Mina’s total receipt in playing the trail game on $\llbracket-j-1,k+1\rrbracket$, the counter stopping on arrival at j1-j-1 or $k+1$ with terminal payments given by $(m_{-j-1},m_{k+1},n_{-j-1},n_{k+1})$.
@@
Let $i\in\mathbb{Z}$ be given, and let $S_{-}^{\mathrm{dev}}\in\mathcal{S}$ be deviating for Mina. Suppose that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}(E)=1$. For any given $(u,\ell)\in\mathsf{D}_{-}(S_{-}^{\mathrm{dev}},i)$,
$\sup\,\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}^{j,k}]\,<\,\mathbb{E}_{S_{-},S_{+}}^{i}[P_{-}]\,$,
@@
Let $i\in\mathbb{Z}$, and let $S_{-}^{\mathrm{dev}}\in\mathcal{S}$ be deviating for Mina. Then
$\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]<\mathbb{E}_{S_{-},S_{+}}^{i}[P_{-}]\,.$
This pair of propositions forms the backbone of the proof of Theorem 1.8(2). They are simply the assertions made by [26, Propositions $4.11$ and $4.12$] in regard to Mina’s deviation. (For a reason to be explained shortly, Proposition 4.16 is phrased a little differently than [26, Proposition $4.11(1)$] and includes a new hypothesis.) Alongside symmetric assertions regarding Maxine’s deviant play made in these results from [26] but omitted here, Theorem 1.8(2) follows directly. Indeed, Mina’s replacement of $S_{-}$ by another strategy $S$ when playing against $S_{+}$ will either effect no change in her mean outcome—namely, $\mathbb{E}^{i}_{S,S_{+}}[P_{-}]=\mathbb{E}^{i}_{S_{-},S_{+}}[P_{-}]$—if $S$ is not deviating; or a negative change, $\mathbb{E}^{i}_{S,S_{+}}[P_{-}]<\mathbb{E}^{i}_{S_{-},S_{+}}[P_{-}]$, by Proposition 4.17. And of course likewise if Maxine is the one to deviate.
@@
Our presentation of the proof of Theorem 1.8(2) is intended to be comprehensive in describing changes to the counterpart in [26, Section $4.2$], and to offer a substantially complete conceptual guide to the proof while avoiding excessive repetition of [26]. We will begin by describing the more major change, which concerns the proof of Proposition 4.17 and will entail presenting a further result, Proposition 4.18. We will describe why this result is needed and state it. An overview of the derivation at large will then be offered, in which some more minor changes to the proof in [26] will be noted. Then we will prove Proposition 4.18.
@@ 4.5.1. The substantial new element, which handles possible non-escape
In the proof of [26, Proposition 4.11], counterpart to Proposition 4.17, the case $\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})>0$ of possible non-escape is treated separately, by a simple argument asserting that, in this case, $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]=$-$\infty$ while [^5] That $\mathbb{E}_{S_{-},S_{+}}^{i}[P_{+}]=m_{i}$ and $\mathbb{E}_{S_{-},S_{+}}^{i}[P_{-}]=n_{i}$ is proved in [26, Lemma $3.11(2)$] which is contingent on [26, Lemma $3.7$]. The latter result has an invalid proof for the present context (where $\kappa$ may be less than one), but in the application in question, the pair $(S_{-},S_{+})$ lies in $\mathcal{S}_{0}^{2}$ with the stake amounts $a_{i}$ and $b_{i}$ all being positive; and, in this case, [26, Lemma $3.7$] is readily obtained for $\kappa\in(0,1)$. $\mathbb{E}_{S_{-},S_{+}}^{i}[P_{-}]=n_{i}>$-$\infty$. The conclusion that $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]=$-$\infty$ is easy to reach in the pure stake $\kappa=1$ case: since $\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})>0$, an edge $[i,i+1]$ indexed by some $i\in\mathbb{Z}$ may be found that is traversed from right-to-left infinitely often with positive probability. When the counter is at $i+1$, Mina consistently faces a stake of $a_{i+1}>0$, so that, in order to win infinitely many of the moves from site $i+1$, she has to an expend infinitely in stake payments. In the present case, where $\kappa\in(0,1)$, this reasoning is flawed, because each move from site $i+1$ is flip with probability $1-\kappa>0$, so that the edge $[i,i+1]$ may in principle be traversed from right to left by the counter on infinitely many occasions without Mina spending a dime when the counter is at $i+1$.
We will circumvent this difficulty: rather than establishing that $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}[P_{-}]=$-$\infty$ when $\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})$ is positive, we will invoke the next result. We write $\mathsf{Trap}$ for the complement of the escape event $E$.
@@
Suppose that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{i}(\mathsf{Trap})>0$. There exists an altered strategy for Mina $S_{-}^{\mathrm{dev}}[\mathrm{alt}]\in\mathcal{S}$ such that
$\textrm{$\mathbb{P}^{i}_{S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}}(\mathsf{Trap}^{c})=0$ and $\mathbb{E}^{i}_{S_{-}^{\mathrm{dev}},S_{+}}[P_{-}]\leq\mathbb{E}^{i}_{S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}}[P_{-}]$}\,.$ (37)

### v4

@@ 4.3. Positive stakes at Nash equilibrium
Derivation. Under $\mathbb{P}^{i}_{S_{-},S_{+}}$, let $\sigma_{i+1}\in\mathbb{N}_{+}\cup\{\infty\}$ denote the stopping time $\inf\big\{\ell\in\mathbb{N}_{+}:X_{\ell}=i+1\big\}$. In the specification (1.3) of Maxine’s net receipt $P_{+}$ as $T_{+}-C_{+}$, the running cost $C_{+}$ may be written $C_{+}\llbracket 1,t\rrbracket$ and $C_{+}\llbracket t+1,\infty\rrbracket$ where Maxine’s stakes up to the $t$th turn enter as summands in the first term. Taking $t=\sigma_{i+1}$, we use the strong Markov property at time $\sigma_{i+x1}$ and drop $C_{+}\llbracket 1,\sigma_{i+1}\rrbracket\geq 0$ to obtain
@@
Here, the left-hand side equals $m_{i}$ by definition while the right-hand side is
@@
The third right-hand term vanishes by Proposition 4.1, so that $m_{i}$ is bounded above by a weighted average of $m_{-\infty}$ and $m_{i+1}$. We will find as desired that $m_{i}\leq m_{i+1}$ by showing $m_{-\infty}\leq m_{i+1}$. In this regard, we first claim that $\mathbb{E}^{i+1}_{S_{-},0}[P_{+}]=m_{-\infty}$. To check this, note that Lemma 4.7(3) implies that $S_{-}$ is wide. We may now make use of Proposition 4.8 to learn that $E_{-}$, and thus also $T_{+}=m_{-\infty}$, are $\mathbb{P}^{i+1}_{S_{-},0}$-almost certain. The absence of running costs for Maxine means that $P_{+}=T_{+}$ under $\mathbb{P}^{i+1}_{S_{-},0}$. The claim obtained, we use it and $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$ to find that $m_{i+1}\,=\,\mathbb{E}^{i+1}_{S_{-},S_{+}}[P_{+}]\,\geq\,\mathbb{E}^{i+1}_{S_{-},0}[P_{+}]=m_{-\infty}$, thereby confirming $m_{i}\leq m_{i+1}$. Omitting the similar proof that $n_{i+1}\leq n_{i}$, we obtain Lemma 4.11. ∎
@@
$m_{i}\,=\,\Big(\kappa\tfrac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i+1}+\Big(\kappa\tfrac{b_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}+\tfrac{1-\kappa}{2}\Big)m_{i-1}\,-\,a_{i}\,.$ (4.7)
Proof. Maxine will spend $a_{i}$ at the first turn; the move will be stake with probability $\kappa$ and then she win it with conditional probability $\tfrac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}$; if she does so, the counter will reach $i+1$, and her resulting conditional mean receipt will be $m_{i+1}$; and this circumstance will equally arise if a fair coin lands heads on a flip move, with probability $(1-\kappa)/2$. Otherwise, Maxine’s receipt will be $m_{i-1}$. Note that the two ratios on the right-hand side of (4.7) are well defined, because $a_{i}+b_{i}>0$. ∎
@@
Proof of Proposition 4.10. By Lemma 4.7(3), $S_{+}$ is wide. By Lemma 4.14, $a_{i}>0$ implies that $a_{i+1}>0$. Hence, all coefficients $a_{i}$ are positive; by Lemma 4.14(2), so are all the $b_{i}$. By Lemma 4.13, the differences $m_{i,i+1}$ and $n_{i+1,i}$ are also found to be positive.∎
@@ 4.4. The forward implication
We are ready for the next derivation. The argument follows the lines of the proof of [24, Theorem $2.6(1)$], with a different approach used at the end to handle flip moves.
@@
$A_{+}-O_{+}\,=\,\bigg(\frac{\rho\,a_{i}^{\rho-1}b_{i}^{\rho}}{(a_{i}^{\rho}+b_{i}^{\rho})^{2}}\kappa\,m_{i-1,i+1}-1\bigg)\cdot\eta\cdot\big(1+o(1)\big)\,$, (4.8)
@@
$\rho\,a_{i}^{\rho}b_{i}^{\rho-1}\,\kappa\,n_{i+1,i-1}\;=\;\big(a_{i}^{\rho}+b_{i}^{\rho}\big)^{2}\,.$ (4.10)
The obtained equations (4.9) and (4.10) are ABMN(3,4)(3,4) with index $i$.
We have shown that $\big\{(a_{i},b_{i},m_{i},n_{i}):i\in\mathbb{Z}\big\}$ is an element of ABMN($\kappa,\rho$). To finish the proof of Theorem 1.3(1), it remains to confirm that the boundary values (1.5) are achieved. We will prove that $\lim_{i\to\infty}m_{-i}=m_{-\infty}$; the three other limits are similarly obtained. The sequence $\big\{m_{-i}:i\in\mathbb{N}\big\}$ decreases by Proposition 4.10 to a limit that we call $\mathfrak{m}_{-\infty}$.
By Definition 1.1 and $(S_{-},S_{+})\in\mathcal{N}_{\kappa,\rho}$, $m_{i}=\mathbb{P}^{i}_{S_{-},S_{+}}[P_{+}]\geq\mathbb{P}^{i}_{S_{-},0}[P_{+}]$. It is Proposition 4.8 that now permits us to identify the right-hand term as being equal to $m_{-\infty}$. Hence, $\mathfrak{m}_{-\infty}\geq m_{-\infty}$; we wish to obtain the opposing inequality. We take the mean of the equality $P_{+}=T_{+}-C_{+}$ in (1.3) and remove non-negative running costs $C_{+}$ to find that $m_{i}\leq\mathbb{P}^{i}_{S_{-},S_{+}}(E_{-})\cdot m_{-\infty}+\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})\cdot m_{\infty}$ where we invoked Proposition 4.1 to eliminate a non-escape $E^{c}$ term. Thus $\mathfrak{m}_{-\infty}\leq m_{-\infty}$ provided that we show that $\lim_{i\to-\infty}\mathbb{P}^{i}_{S_{-},S_{+}}(E_{+})$ equals zero: far to the left, Mina’s victory is close to assured.
@@ 4.5. The reverse implication
Here we prove Theorem 1.3(2), the step at which the infinite-turn game is controlled by comparison with finite-trail counterparts. Throughout, $\big\{(a_{i},b_{i},m_{i},n_{i}):i\in\mathbb{Z}\big\}$ denotes an element of ABMN($\kappa,\rho$), with boundary data $(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})$ that satisfies (1.6). We define strategies $S_{-},S_{+}\in\mathcal{S}$ that offer $b$- and $a$-stake compatibly with the rule (1.4).
Since all counter moves are $\pm 1$, counter location is constrained by parity. First we denote the set of space-time sites that are thus in principle accessible for gameplay $X:\mathbb{N}\to\mathbb{Z}$ under $\mathbb{P}^{h}_{S_{1},S_{2}}$ for some strategy pair $(S_{1},S_{2})\in\mathcal{S}^{2}$. Here and throughout the section, the symbol $h$ is reserved for denoting a choice of initial counter location.
@@
For $h\in\mathbb{Z}$, the forward play-cone $F_{h}$ of $h$ is set equal to
$F_{h}\,=\,\Big\{\,(k,\ell)\in\mathbb{Z}\times\mathbb{N}_{+}:|k-h|\leq\ell\,,\,|k-h|+\ell\in 2\mathbb{N}\,\Big\}\,.$
Let $S\in\mathcal{S}$ (and recall the formulation of the strategy space $\mathcal{S}$ from Section 1.3.1). A Mina deviation point is an element $(q,\ell)\in F_{h}$ for which there exists a trajectory $\psi:\llbracket 0,\ell\rrbracket\to\mathbb{Z}$ with $\psi(0)=h$ and $\psi(\ell)=q$ such that $S(\psi)\neq b_{q}$. Write $\mathsf{D}_{-}(S,h)\subseteq F_{h}$ for the set of Mina deviation points. The strategy $S$ is deviating for Mina if $\mathsf{D}_{-}(S,h)\neq\emptyset$. A Maxine deviation point is an element $(q,\ell)\in F_{h}$ such that $S(\psi)\not=a_{q}$ for some path $\psi$ as above. Write $\mathsf{D}_{+}(S,h)$ for the set of these points; if $\mathsf{D}_{+}(S,h)\not$$\emptyset$, then $S$ is deviating for Maxine.
@@
Write $P_{-}^{j,k}$ for Mina’s total receipt in playing the trail game on $\llbracket-j-1,k+1\rrbracket$, the counter stopping on arrival at j1-j-1 or $k+1$ with terminal payments given by $(m_{-j-1},m_{k+1},n_{-j-1},n_{k+1})$.
@@
Let $h\in\mathbb{Z}$ be given, and let $S_{-}^{\mathrm{dev}}\in\mathcal{S}$ be deviating for Mina. Suppose that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}(E)=1$. For any given $(u,\ell)\in\mathsf{D}_{-}(S_{-}^{\mathrm{dev}},h)$,
$\sup\,\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}^{j,k}]\,<\,\mathbb{E}_{S_{-},S_{+}}^{h}[P_{-}]\,$,
@@
Let $h\in\mathbb{Z}$, and let $S_{-}^{\mathrm{dev}}\in\mathcal{S}$ be deviating for Mina. Then
$\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]<\mathbb{E}_{S_{-},S_{+}}^{h}[P_{-}]\,.$
This pair of propositions forms the backbone of the proof of Theorem 1.3(2). They are simply the assertions made by [24, Propositions $4.11$ and $4.12$] in regard to Mina’s deviation. (For a reason to be explained shortly, Proposition 4.16 is phrased a little differently than [24, Proposition $4.11(1)$] and includes a new hypothesis.) Alongside symmetric assertions regarding Maxine’s deviant play made in these results from [24] but omitted here, Theorem 1.3(2) follows directly. Indeed, Mina’s replacement of $S_{-}$ by another strategy $S$ when playing against $S_{+}$ will either affect no change in her mean outcome—namely, $\mathbb{E}^{h}_{S,S_{+}}[P_{-}]=\mathbb{E}^{h}_{S_{-},S_{+}}[P_{-}]$—if $S$ is not deviating; or a negative change, $\mathbb{E}^{h}_{S,S_{+}}[P_{-}]<\mathbb{E}^{h}_{S_{-},S_{+}}[P_{-}]$, by Proposition 4.17. And of course likewise if Maxine is the one to deviate.
@@
Our presentation of the proof of Theorem 1.3(2) is intended to be comprehensive in describing changes to the counterpart in [24, Section $4.2$], and to offer a substantially complete conceptual guide to the proof while avoiding excessive repetition of [24]. We will begin by describing the more major change, which concerns the proof of Proposition 4.17 and will entail presenting a further result, Proposition 4.18. We will describe why this result is needed and state it. An overview of the derivation at large will then be offered, in which some more minor changes to the proof in [24] will be noted. Then we will prove Proposition 4.18.
@@ 4.5.1. The substantial new element, which handles possible non-escape
In the proof of [24, Proposition 4.11], counterpart to Proposition 4.17, the case $\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})>0$ of possible non-escape is treated separately, by a simple argument asserting that, in this case, $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]=$-$\infty$ while [^3] That $\mathbb{E}_{S_{-},S_{+}}^{h}[P_{+}]=m_{h}$ and $\mathbb{E}_{S_{-},S_{+}}^{h}[P_{-}]=n_{h}$ is proved in [24, Lemma $3.11(2)$] which is contingent on [24, Lemma $3.7$]. The latter result has an invalid proof for the present context (where $\kappa$ may be less than one), but in the application in question, the pair $(S_{-},S_{+})$ lies in $\mathcal{S}_{0}^{2}$ with the stake amounts $a_{i}$ and $b_{i}$ all being positive; and, in this case, [24, Lemma $3.7$] is readily obtained for $\kappa\in(0,1)$. $\mathbb{E}_{S_{-},S_{+}}^{h}[P_{-}]=n_{h}>$-$\infty$. The conclusion that $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]=$-$\infty$ is easy to reach in the pure stake $\kappa=1$ case: since $\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})>0$, an edge $[i,i+1]$ indexed by some $i\in\mathbb{Z}$ may be found that is traversed from right-to-left infinitely often with positive probability. When the counter is at $i+1$, Mina consistently faces a stake of $a_{i+1}>0$, so that, in order to win infinitely many of the moves from site $i+1$, she has to expend infinitely in stake payments. In the present case, where $\kappa\in(0,1)$, this reasoning is flawed, because each move from site $i+1$ is flip with probability $1-\kappa>0$, so that the edge $[i,i+1]$ may in principle be traversed from right to left by the counter on infinitely many occasions without Mina spending a dime when the counter is at $i+1$.
We will circumvent this difficulty: rather than establishing that $\mathbb{E}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}[P_{-}]=$-$\infty$ when $\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}(E^{c})$ is positive, we will invoke the next result. We write $\mathsf{Trap}$ for the complement of the escape event $E$.
@@
Suppose that $\mathbb{P}_{S_{-}^{\mathrm{dev}},S_{+}}^{h}(\mathsf{Trap})>0$. There exists an altered strategy for Mina $S_{-}^{\mathrm{dev}}[\mathrm{alt}]\in\mathcal{S}$ such that
$\textrm{$\mathbb{P}^{h}_{S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}}(\mathsf{Trap})=0$ and $\mathbb{E}^{h}_{S_{-}^{\mathrm{dev}},S_{+}}[P_{-}]\leq\mathbb{E}^{h}_{S_{-}^{\mathrm{dev}}[\mathrm{alt}],S_{+}}[P_{-}]$}\,.$ (4.11)